# DWGS: Enhancing Sparse-View Gaussian Splatting with Hybrid-Loss Depth Estimation and Bidirectional Warping

### v1

@@ IV-B5 Efficiency
The proposed DWGS method matches the efficiency of other advanced approaches, achieving an inference speed of 250 FPS, which is comparable to methods such as SCGaussian and MCGaussian, while maintaining a low average training time of 2.5 minutes. In terms of GPU memory usage, DWGS optimizes resource consumption, requiring 3 GB for the Hybrid-Loss Depth Estimation (HLDE) component and 5 GB for the Virtual View Synthesis (VVS) component, outperforming methods like FreeNeRF (4×48 GB) and SparseNeRF (32 GB). Compared to approaches such as 3DGS (7.5 minutes, 2 GB) and FSGS (17 minutes, 3 GB), DWGS offers a substantial reduction in training time while preserving real-time inference capabilities, making it a highly efficient solution for practical 3D reconstruction tasks.
TABLE VI: Efficiency comparison of different methods.
@@
FreeNeRF[8] $9\times 10^{-2}$ 2.3 h $\times4$8 GB
SparseNeRF[7] $9\times 10^{-2}$ 1.5 h 32 GB
3DGS[5] 400 7.5 min 2 GB
FSGS[26] 300 17 min 3 GB
SID[26] 300 28 min 3.5 GB
DNGaussian[27] 500 3.5 min 2 GB
MCGS[36] 250 2.5 min 2 GB
SCGaussian[37] 250 1.5 min 3 GB
DWGS 250 2.5 min
@@ 5 GB (VVS)
IV-C Ablation
An ablation study is conducted on the LLFF dataset under a 3-view setting to evaluate the individual contributions of the core components in DWGS: Hybrid-Loss Depth Estimation (HLDE), Bidirectional Warping Virtual View Synthesis (VVS), and Occlusion-Aware Reconstruction (OAR). Quantitative results are summarized in Table VII. The baseline methods, vanilla 3DGS and SCGaussian, achieve a PSNR of 15.42 dB and 20.73 dB, respectively, with LPIPS scores of 0.463 and 0.196. The combination of HLDE + VVS + OAR (full DWGS) yields the best performance, achieving a PSNR of 21.13 dB and an LPIPS of 0.189.
Figure 13 provides a qualitative comparison of different component configurations on the Fern scene. In the boxed regions, the following observations are made: +HLDE: Produces more robust depth estimates, effectively resolving structural distortions present in 3DGS. +VVS: Introduces virtual view constraints, significantly reducing artifacts and improving consistency. +OAR: Generates more plausible background content and alleviates occlusion-related artifacts.
Compared to the baseline methods, DWGS produces renderings that are visually closer to the ground truth, demonstrating the effectiveness of the proposed components.
+HLDE        +HLDE+VVS     +HLDE+VVS+OAR
depth map        3DGS          GT
Figure 13: Ablation study on component contributions. Compared to 3DGS, the rendering results of full DWGS model are closest to GT.
TABLE VII: Ablation study on the components of DWGS.
@@ V Conclusion
Limitations: DWGS relies on COLMAP for pose estimation and uses fixed thresholds in its propagation constraints. The occlusion handling module is currently optimized for foreground content and may not generalize well to sky or homogeneous regions. DWGS is tailored for static scenes and does not support dynamic content.
Conclusion: This paper presents DWGS, an effective framework for sparse-view NVS based on 3DGS. The core of our approach includes three innovations: a Hybrid-Loss Depth Estimation module that enforces multi-view consistency through reprojection and point propagation constraints; a Bidirectional Warping Virtual View Synthesis method that enhances the coverage of unobserved regions, thereby mitigating overfitting to the limited input views; and an Occlusion-Aware Reconstruction component that improves background rendering through depth-difference priors. Extensive experiments across standard benchmarks demonstrate that DWGS achieves state-of-the-art performance under extreme sparsity. We believe our work offers a valuable step toward practical and high-quality 3D reconstruction from very few images.

### v2

@@ IV-B5 Efficiency
Efficiency evaluation of different methods on the LLFF dataset in Table VII. The proposed HBSplat method matches the efficiency of other advanced approaches, achieving an inference speed of 250 FPS, which is comparable to methods such as SCGaussian and MCGaussian, while maintaining a low average training time of 2.5 minutes. In terms of GPU memory usage, HBSplat optimizes resource consumption, requiring 3 GB for the Hybrid-Loss Depth Estimation (HLDE) component and 5 GB for the Virtual View Synthesis (VVS) component, outperforming methods like FreeNeRF (4×48 GB) and SparseNeRF (32 GB). Compared to approaches such as 3DGS (7.5 minutes, 2 GB) and FSGS (17 minutes, 3 GB), HBSplat offers a substantial reduction in training time while preserving real-time inference capabilities, making it a highly efficient solution for practical 3D reconstruction tasks.
TABLE VII: Efficiency comparison of different methods.
@@
FreeNeRF
[CVPR23]
[8]
$9\times 10^{-2}$ 2.3 h $\times4$8 GB
SparseNeRF
[ICCV23]
[7]
$9\times 10^{-2}$ 1.5 h 32 GB
3DGS
[SIGGRAPH23]
[5]
400 7.5 min 2 GB
FSGS
[ECCV24]
[26]
300 17 min 3 GB
SID
[ICASSP25]
[26]
300 28 min 3.5 GB
DNGaussian
[CVPR24]
[27]
500 3.5 min 2 GB
MCGS
[TPAMI25]
[36]
250 2.5 min 2 GB
SCGaussian
[NeurIPS24]
[37]
250 1.5 min 3 GB
HBSplat(Ours) 250 2.5 min
@@ 5 GB (VVS)
IV-C Ablation Study
Ablation study is conducted on the LLFF dataset under a 3-view setting to evaluate the individual contributions of the core components in HBSplat: Hybrid-Loss Depth Estimation (HLDE), Bidirectional Warping Virtual View Synthesis (VVS), and Occlusion-Aware Reconstruction (OAR).
Quantitative Component Analysis: Quantitative results are summarized in Table VIII. The full HBSplat model (HLDE+VVS+OAR) yields the best overall performance, achieving a PSNR of 21.13 dB and an LPIPS of 0.189. Among the components, VVS contributes the highest PSNR due to the strong photometric constraints imposed by virtual views. Although HLDE yields a lower PSNR than VVS, it infers more accurate geometry, resulting in the lowest LPIPS. Notably, each individual component of HBSplat outperforms all other baselines. Furthermore, the performance gain from combining components is non-linear, integrating HLDE with VVS alone is sufficient to achieve substantial improvement.
TABLE VIII: Ablation study on the components of HBSplat.
@@ IV-C Ablation Study
Qualitative Component Analysis: Figure 13 provides a qualitative comparison on the Fern scene. The following observations are made in the boxed regions: +HLDE: Produces more robust depth estimates, effectively resolving structural distortions present in 3DGS. +VVS: Introduces virtual view constraints, significantly reducing artifacts and improving consistency. +OAR: Generates more plausible background content and alleviates occlusion-related artifacts.
In the second row of the comparison, SCGaussian exhibits black artifacts and fails to reconstruct the background completely, while MCGS reconstructs leaves as sparse, needle-like structures and suffers from blurred edges. In contrast, HBSplat produces renderings that are visually closer to the GT, demonstrating the effectiveness of the proposed components.
PPC with different nearest neighbor distance ($d_{NN}$): Figure 14 shows the line graph of PSNR values over different $d_{NN}$. It can be observed that as $d_{NN}$ increases, both the number of common points and the PSNR value gradually rise. However, when the $d_{NN}$ exceeds 3, the PSNR for both the Fern scene and the LLFF dataset begins to decline. The point (22.38, 99) represents the PSNR value and the number of common points for the Fern scene, respectively. Notably, if the number of common points is less than 100 following a secondary filtering step, the PPC operation is skipped to avoid redundant processing.
+HLDE        +HLDE+VVS     +HLDE+VVS+OAR
MCGS       SCGaussian         GT
Figure 13: Ablation study on component contributions. Compared to MCGS and SCGaussian, the rendering results of full HBSplat model are closest to.
Figure 14: PSNR vs. $d_{NN}$. Both the number of common points and PSNR initially increase with $d_{NN}$ but decline after the $d_{NN}$ exceeds 3.
@@ V Conclusion
This paper presents HBSplat, an effective framework for sparse-view NVS based on 3DGS. The core of our approach includes three innovations: a Hybrid-Loss Depth Estimation module that enforces multi-view consistency through reprojection, point propagation, and smoothness constraints; a Bidirectional Warping Virtual View Synthesis method that enhances the coverage of unobserved regions, thereby mitigating overfitting to the limited input views; and an Occlusion-Aware Reconstruction component that improves background rendering through depth-difference priors. Extensive experiments across standard benchmarks demonstrate that HBSplat achieves state-of-the-art performance under extreme sparsity. We believe our work offers a valuable step toward practical and high-quality 3D reconstruction from very few images.
Limitations and Future Work: HBSplat has several limitations. It relies on COLMAP for pose estimation and uses fixed propagation thresholds, potentially limiting adaptability. The occlusion module, optimized for foreground objects, may struggle with large homogeneous regions such as sky. Additionally, it only supports static scenes. Future work will explore self-supervised pose estimation, adaptive thresholds, and dynamic scene modeling to address these constraints.